# NeuCo-Bench: A Novel Benchmark Framework for Neural Embeddings in Earth Observation

### v1

@@ B.3 General Evaluation Setup
B.3.1 Method Details
Averaging Baseline. As a simple informative reference, we construct a Mean baseline by strongly downsampling and averaging the SSL4EO-S12 data cubes. First, we reduce the spatial resolution of each of the 27 channels from 264x264 pixels to 8x8 by bi-linear interpolation. Next, we exploit correlation as visualized by Fig.˜10 reducing the number of channels from 27 down to four. We average the channels B1 through B9 of both S2L1C and S2L2A and we do similar for channels B11 and B12, respectively. Channel B10 of S2L1C is kept separate since no corresponding band exists in S2L2A. Seasonal snapshots are kept separate, yielding $\times8$$\times$$\times4$ values, flattened to $N=1024$. This baseline indicates how much task-relevant information survives coarse spatial and spectral aggregation.
@@ B.3.1 Method Details
Neural rate–distortion compressors. We adopt the Factorized Prior autoencoder of Ballé et al. (2016) for Sentinel-2 L1C imagery. Models use 256 intermediate channels and 128 latent channels, trained with loss $\mathcal{L}=R+\lambda D$, where $D$ is MSE distortion and $\lambda\in\{0.025,0.1,0.5\}$. These compressors match or exceed JPEG2000 PSNR at roughly half the bitrate (Fig. 11). At inference, we pool latents to $4\times 4$, flatten to 2048, and average adjacent channels to yield 1024-dim embeddings.
Self-supervised foundation models (FMs). We evaluate unimodal FMs pretrained on Sentinel-2 L1C (Wang et al., 2023b), as well as multimodal FMs (DOFA (Xiong et al., 2024), TerraMind (Jakubik et al., 2025)). CNN backbones are downsampled by spatial averaging to 2048 dims, then halved to 1024. ViT-B/16 backbones average patch embeddings, padding from 768 to 1024 dims.
Temporal aggregation. Each sample contains four seasonal snapshots. In pre-encoding, channels are averaged across seasons before encoding. In post-encoding, each snapshot is encoded separately and averaged. As shown in Fig. 9, post-encoding better handles seasonal outliers (e.g., snow), at the cost of increased compute.
B.3.2 Key Results Across Embedding Methods
Table˜4 summarizes linear-probe $R^{2}$ scores with pre- and post-encoding aggregation.
Temporal aggregation. Post‐encoding aggregation yields absolute $R^{2}$ gains for almost every method and task. For instance, the TerraMind cloud performance increases from 0.434 to 0.731 (+0.297), and landcover tasks by +0.011–0.016.
Neural compressors. Reducing $\lambda$ from 0.1 to 0.025 yields small average improvements, whereas $\lambda=0.5$ degrades performance and produces negative $R^{2}$ on several tasks.
ViT backbones. MAE achieves balanced performance across tasks, while contrastive DINO and MoCo excel on semantic landcover tasks (e.g., DINO post‐encoding achieves $R^{2}=0.863$ on crop fraction) but are less competitive on geophysical regression.
CNN backbones. Although CNN variants perform strong on landcover tasks, they fall behind ViTs on geophysical regressions. They benefit from post‐encoding aggregation—for instance, DINO CNN’s biomass mean improves from negative scores to 0.117.
Multimodal models. TerraMind consistently outperforms all other embeddings, achieving the highest $R^{2}$ across tasks. DOFA, while generally scoring below TerraMind, still achieves consistent performance across all task.
Table 4: Full per‐task $R^{2}$ scores for all tested embedding methods. For all methods, we report scores for Pre-Encoding Temporal Aggregation (Pre-Enc) and Post-Encoding Temporal Aggregation (Post-Enc). Methods are grouped into the main categories: Multimodal (Terramind, DOFA), ViT FMs, CNN FMs, Neural Compressors, and Averaging Baseline.
Method Biomass Mean Biomass Std Crops Clouds Landcover Agriculture Landcover Forest Heatisland Mean Heatisland Std
Averaging Baseline -0.552 -0.426 -0.061 -2.293 0.126 0.216 -0.962 -1.157
Neural Compressors (Factorized Prior)
$\lambda$ = 0.025 (Pre-Enc) 0.186 0.132 0.423 0.244 0.463 0.246 0.321 -0.114
$\lambda$ = 0.025 (Post-Enc) 0.195 0.140 0.338 0.288 0.449 0.256 0.315 -0.113
$\lambda$ = 0.1 (Pre-Enc) 0.127 0.078 0.388 0.222 0.447 0.278 0.238 -0.246
$\lambda$ = 0.1 (Post-Enc) 0.129 0.078 0.357 0.266 0.464 0.333 0.219 -0.270
$\lambda$ = 0.5 (Pre-Enc) -0.037 -0.062 0.364 0.129 0.464 0.319 0.067 -0.563
$\lambda$ = 0.5 (Post-Enc) -0.036 -0.071 0.325 0.200 0.478 0.343 0.033 -0.628
CNN FMs (Unimodal)
DINO (Pre-Enc) -0.031 0.015 0.749 -0.171 0.855 0.838 0.396 -0.334
DINO (Post-Enc) 0.117 0.088 0.826 0.147 0.879 0.865 0.483 -0.221
MoCo (Pre-Enc) -0.164 -0.111 0.688 -0.228 0.835 0.808 0.236 -0.502
MoCo (Post-Enc) -0.048 -0.030 0.780 0.216 0.864 0.844 0.345 -0.334
ViT FMs (Unimodal)
DINO (Pre-Enc) 0.159 0.157 0.755 0.061 0.822 0.812 0.446 -0.169
MoCo (Pre-Enc) 0.294 0.229 0.706 0.242 0.790 0.778 0.477 0.113
MoCo (Post-Enc) 0.338 0.259 0.751 0.409 0.825 0.814 0.506 0.133
Multimodal FMs
TerraMind (Pre-Enc) 0.492 0.366 0.836 0.434 0.907 0.892 0.676 0.193
TerraMind (Post-Enc) 0.528 0.390 0.879 0.731 0.918 0.908 0.691 0.226
DOFA (Pre-Enc) 0.311 0.234 0.622 0.409 0.762 0.747 0.521 0.156
DOFA (Post-Enc) 0.373 0.269 0.587 0.518 0.777 0.773 0.560 0.207
Intermediate Layers. To probe how different layers capture task‐relevant features, we extracted intermediate embeddings from the contrastive ViTs. As shown in Fig.˜12, low‐level regression tasks (biomass mean/std, heat‐island mean/std, cloud fraction) peak at shallow layers (layer 1), whereas higher‐order semantic tasks (crop fraction, landcover proportion) reach their best performance at deeper layers.
Figure 12: Layer-wise $R^{2}$ performance of ViT-based foundation models across downstream tasks. Shallow layers capture low-level signals useful for regression tasks (e.g., biomass, heat-islands, clouds), while deeper layers improve performance on semantic classification tasks (e.g., crops, landcover).
B.3.3 Ablations and Extended Analyses
Decoder Analysis (Linear vs. Non-Linear Probes).
@@ B.3.3 Ablations and Extended Analyses
Figures 5(b) and 13 compare linear probes with one- and two-hidden-layer MLP decoders. We observe three consistent findings:
@@
Figures 14 and 15 show ablation results on embedding dimensionality for ViT-based and CNN-based models, respectively. For CNN backbones, performance generally peaks in the range of 128–1024 dimensions, with larger or smaller embeddings leading to consistent performance drops. ViT-based embeddings, by contrast, are most effective at their natural patch-token dimension, and reductions tend to degrade task performance. Notably, the benefit of larger embeddings is limited: increases beyond 1024 dimensions yield negligible accuracy improvements while substantially raising computational demands and probe parameter counts. These results justify the use of 1024-dimensional embeddings as a balanced default in the run data challenge, while also illustrating NeuCo-Bench’s flexibility for exploring embedding-size vs. utility trade-offs in future studies.
B.4 Compute Resources & Tuning Parameters
Task (# samples) $N=1024$ $N=512$ $N=2048$
Biomass (2415) 4.23 4.24 7.98 16.01 16.54 3.87 7.54 4.25 8.32
Crops (3355) 5.60 5.56 11.26 22.15 22.00 5.09 10.35 5.94 12.34
Clouds (1140) 2.04 2.03 3.94 7.76 7.59 1.91 3.69 2.13 4.12
Landcover Forest (4691) 7.86 7.81 15.48 31.06 30.47 7.08 14.37 8.06 16.64
Heatisland (1659) 2.79 2.80 5.53 10.84 10.83 2.66 5.25 2.89 5.85
No-data (13260) 22.45 22.45 44.30 88.68 88.18 20.46 42.03 23.18 46.60
  • Embedding dimension ($N$, embedding_dim)
  • Number of epochs per CV fold ($E$, epochs)
  • Number of CV folds ($k$, k_folds)
  • Choice of tasks included (task_filter)
Empirical runtime measurements confirm an approximately linear scaling w.r.t. both, the number of epochs $E$ and the number of cross-validation folds $k$. A similar scaling behavior was numerically verified for the dataset size (# samples) for fixed downstream task. Runtimes are further influenced–—though to lesser extent—–by the embedding dimensionality $N$. For example, increasing the embedding size from 512 to 1024 dimensions results in a runtime increase of approximately 5% to 10% across tasks. For $N=1024$ to $N=2048$ dimensions implies an additional increase of about 5% to 15%—depending on the task dataset size. Such a sub-linear scaling may be attributed to computation overheads and system-level inefficiencies. Those may reduce the relative computational costs when increasing the embedding dimensionality $N$. Table˜5 lists a collection of recorded execution times (in seconds) for various parameter configurations per downstream task. All runtimes were measured on a single commodity ARM64 CPU with 16 cores (4.06 GHz) and 64 GB of RAM.
B.5 Licenses for Data & Software
NeuCo-Bench builds on open-source software and is released under the Apache 2.0 license publicly available at https://github.com/embed2scale/NeuCo-Bench. All package dependencies are listed in the requirements.txt [^25] https://github.com/embed2scale/NeuCo-Bench/blob/main/requirements.txt file, and those are licensed under widely-accepted open-source terms [^26] https://opensource.org/licenses, including BSD, MIT, PSFL, The Unlicense272727code-equivalent to CC0 data licenses, MPL-2.0282828weak copy-left that allows for integration with non-copyleft licenses, and Apache. These permissive licenses allow academic research and commercial use, making them fully compatible with the chosen Apache 2.0 license. Table˜6 lists all data currently included in NeuCo-Bench along with their origin. Google Earth Engine (GEE) (GEE, ) was utilized as the primary platform for downloading downstream task data as introduced in Section˜B.1. Future data and code contributions to NeuCo-Bench are required to be licensed under CC-BY 4.0 and Apache 2.0, respectively.
Table 6: List of licenses related to datasets currently included in our benchmark. All of these, except Clouds, are available in GEE. Table˜1 lists years of target labels, ranging from 2018 through 2024.
Dataset Origin of Data License
Sentinel-1 & -2 ESA / Copernicus CC BY-SA 3.0 IGO
Landsat-8 USGS (Observation and Center, 2020) Public Domain
CDL USDA NASS Cropland Data Layers (Boryan et al., 2011) Public Domain
Copernicus Land Monitoring Service (European Environment Agency (2018), EEA) and Free Access
CloudSen12+ CloudSEN12 project (Aybar et al., 2024) CC0 1.0
GEDI NASA (Dubayah et al., 2022) Public Domain
We note that the current NeuCo-Bench implementation lists CUDA packages covered by a proprietary NVIDIA license [^29] https://docs.nvidia.com/cuda/eula/index.html. However, we do neither bundle nor redistributes corresponding binaries. Users and contributors to NeuCo-Bench that share related docker containers need to explicitly attribute NVIDIA’s license. Fortunately, and as alluded in Table˜5 and Section˜A.2, NeuCo-Bench runs swiftly in a VM with commodity hardware specifications on CPU compute, only. Accordingly, the standalone NeuCo-Bench implementation introduced in Section˜B.2 can be started with (Bash) environment variable CUDA_VISIBLE_DEVICES=’’ to avoid usage of GPU resources.

### v2

@@ C.1 Temporal Aggregation Analysis
We study image-based encoding methods for our baseline evaluations. As each input sample contains four seasonal Sentinel-1/2 snapshots, we handle temporal sequence data by reducing the four snapshots into a single embedding using two aggregation strategies:
  • Pre-encoding: seasonal snapshots are averaged before encoding.
  • Post-encoding: each snapshot is encoded separately and the resulting embeddings are averaged. As shown in Fig.˜10, this approach better handles seasonal outliers (e.g., snow) at the cost of additional computation.
Across all tested methods, post-encoding aggregation provides consistent $R^{2}$ improvements for most methods and tasks as summarized in Table˜7. The largest gains occur for the cloud-fraction prediction task, with increases of up to +0.42+0.42 for ViT-based encoders and +0.30+0.30 for TerraMind. Semantic tasks such as land-cover show smaller but consistent improvements (+0.011 to +0.016). In summary, post-encoding yields performance gains, in particular for temporally sensitive tasks such as determination of cloud fractions. These results motivate the use of post-encoding aggregation for the baseline results.
Table 7: Comparison of temporal aggregation methods. We show average $\bar{R}_{p}^{2}$ and $\bar{R}_{P}^{2}$ scores across all downstream tasks (see main paper for details) for pre-encoding vs. post-encoding aggregation, respectively. We also provide the overall improvement $\bm{\Delta R^{2}}=\bar{R}_{P}^{2}-\bar{R}_{p}^{2}$ and the best gain per task, Best $\bm{\Delta R^{2}}$.
Method Pre-Enc. $\bar{R}_{p}^{2}$ Post-Enc. $\bar{R}_{P}^{2}$ $\bm{\Delta R^{2}}$ Best $\bm{\Delta R^{2}}$ (task)
Averaging Baseline -0.522 -0.385 +0.137 +0.270 (clouds)
Factorized Prior 0.238 0.233 -0.005 +0.044 (clouds)
DINO ResNet 0.289 0.397 +0.108 +0.318 (clouds)
DINO ViT 0.382 0.470 +0.088 +0.423 (clouds)
MAE ViT 0.481 0.537 +0.056 +0.272 (clouds)
TerraMind 0.600 0.659 +0.059 +0.297 (clouds)
C.2 Embedding Method Evaluations
The following section extends the baseline evaluations in the main paper and provides additional details on the tested embedding methods.
Averaging Baseline. As a simple informative reference, we construct a Mean baseline by strongly downsampling and averaging the SSL4EO-S12 data cubes. First, we reduce the spatial resolution of each of the 27 channels from 264x264 pixels to 8x8 by bi-linear interpolation. Next, we exploit correlation as visualized by Fig.˜12 reducing the number of channels from 27 down to four. We average the channels B1 through B9 of both S2L1C and S2L2A and we do similar for channels B11 and B12, respectively. Channel B10 of S2L1C is kept separate since no corresponding band exists in S2L2A. Seasonal snapshots are kept separate, yielding $\times8$$\times$$\times4$ values, flattened to $N=1024$. This baseline indicates how much task-relevant information survives coarse spatial and spectral aggregation.
Neural rate–distortion compressors.
We adopt and train a Factorized Prior autoencoder Ballé et al. [5] for Sentinel-2 L1C imagery. Models use 256 intermediate channels and 128 latent channels, trained with loss
$\mathcal{L}=R+\lambda D$ (9)
where $D$ is MSE distortion and $\lambda\in\{0.025,0.1,0.5\}$ and $R$ represents the entropy-coding term for bit-stream compression. These compressors exceed JPEG2000 PSNR at roughly half the bitrate (Fig. 11). At inference, we pool latents to $4\times 4$, flatten to 2048, and average adjacent channels to yield 1024-dim embeddings. We evaluate three rate–distortion settings, which control the weight of the entropy loss term during training, cf. Eq.˜9. We observe that the model with the strongest emphasis on compression ($\lambda=0.025$) performs best overall, while lower compression focus ($\lambda=0.5$) degrades performance. However, all variants score behind FM embeddings and struggle with the strong spatial–temporal aggregation and linear-probing setup.
@@ C.2 Embedding Method Evaluations
Self-supervised foundation models (FMs). We evaluate a broad set of pretrained EO FMs. As introduced in the main paper, we apply a consistent embedding aggregation pipeline across all FMs. As motivated by Section˜C.1, we utilize temporal post-encoding aggregation throughout our experiments. CNN outputs are reduced via global average pooling. If the pooled feature dimension exceeds 1,024, we apply pairwise channel means, i.e., adjacent feature channels are averaged in pairs to halve the dimensionality while preserving coarse channel structure. For ViT encoders, we average the spatial patch tokens (excluding the CLS token, if present) to form a single embedding. Additionally, we also evaluate CLS-token embeddings for Prithvi and Clay. All aggregated embeddings are padded to the target 1,024-dimensional space. Overall, we observe the following trends:
  • FM Embeddings. Multimodal models, most prominently TerraMind, consistently outperforms all other embeddings, achieving the highest $R^{2}$ across most tasks. We highlight that jointly modelling Sentinel-1/2 can provide task benefits under strong spatial and spectral aggregation. DOFA, while scoring below TerraMind, still achieves consistent performance across all task. EO-specific, single-modal ViTs such as Prithvi and Clay follow below TerraMind and offer strong results across semantic and geophysical tasks. SSL4EO-pretrained models exhibit complementary strengths, that is: CNN variants perform strongly on land-cover tasks, but fall behind ViTs on geophysical regressions. Notably, S12-DINO ResNet scores the second highest on both the land-cover agriculture and forest tasks. MAE ViT achieves balanced and high performance across tasks, while contrastive DINO and MoCo excel on semantic land-cover tasks. However, DINO/MoCo are less competitive on geophysical regression.
  • CLS Token Embeddings. The comparison between patch-averaged and CLS-token embeddings for Prithvi and Clay demonstrates that mean patch-token averaging is the more robust strategy: for Prithvi, CLS performance is slightly lower but remains close, whereas for Clay the CLS token underperforms patch averaging.
C.3 Non-linear Probing & Embedding Size Ablations
Linear vs. Non-Linear Probing.
@@ C.3 Non-linear Probing & Embedding Size Ablations
Additionally to the main paper plot on average task scores, we provide per task plots in Figures 13 comparing linear probes with one- and two-hidden-layer MLP decoders. We observe three consistent findings:
@@
Figures 14 and 15 show per-task ablation results on embedding dimensionality for ViT-based and CNN-based models, respectively. For CNN backbones, performance generally peaks in the range of 128–1024 dimensions, with larger or smaller embeddings leading to consistent performance drops. ViT-based embeddings, by contrast, are most effective at their natural patch-token dimension, and reductions tend to degrade task performance. Notably, the benefit of larger embeddings is limited: increases beyond 1024 dimensions yield negligible accuracy improvements while substantially raising computational demands and probe parameter counts. These results justify the use of 1024-dimensional embeddings as a balanced default in the run data challenge, while also illustrating NeuCo-Bench’s flexibility for exploring embedding-size vs. utility trade-offs in future studies.